OpenStudy (anonymous):

what is the area of a regular, NON-right angled triangle with sides 240, 180, and 140 m?

OpenStudy (a_clan):

why is it regular?

OpenStudy (angela210793):

Use Heron's formula:A=\[\sqrt{p(p-a)(p-b)(p-c)}\]

OpenStudy (angela210793):

where p=Perimeter/2 and a,b,c are the lengths if the sides

OpenStudy (anonymous):

Area of a triangle is \[1/2ab \sin C\] Where a b and c are sides and A B and C are the angles opposite to the sides.

OpenStudy (anonymous):

so that is the angle inbetween the two sides, right?

OpenStudy (angela210793):

Heron's formula works always if u have the 3 lengths....why r u mixing it with sin?

OpenStudy (anonymous):

trigonometry. never heard of heron's formula :P but thanks guys :)

OpenStudy (angela210793):

K..here it is http://en.wikipedia.org/wiki/Heron's_formula

OpenStudy (anonymous):

@angela triangles and trigonometry go together like butter and bread! hahaha

OpenStudy (angela210793):

yea i know....but ..am i blind or there's no angle given...

OpenStudy (anonymous):

You use the cosine rule: \[c ^{2} = a ^{2} + b ^{2} - 2ab \cos C\] The reason i used this is that these fomulae are the most commonly used ones. Heron's would be easier but is lesser known and is not usually given on formulae sheets

OpenStudy (anonymous):

@ angela: but you can find that out using the cos rule. CosA= (b^2 + c^2 - a^2)/2bc. I am doing a past paper, so i was asking according to the instructions given.

OpenStudy (angela210793):

I didn't know that...In Albania we use it a looooooooooot...I thought you knew it :)

OpenStudy (anonymous):

ohh. cool B)

OpenStudy (anonymous):

im gonna learn herons, it looks very useful.

OpenStudy (angela210793):

:)lol

OpenStudy (anonymous):

@angela: wait. are you in school/college, what is the "p" in the formula?

OpenStudy (angela210793):

finished hs and I've written above that p=P/2( half of perimeter )

OpenStudy (angela210793):

read the article i sent cause you'll get it

OpenStudy (anonymous):

sry, sry, didn't see that there :P

OpenStudy (angela210793):

:)it's k